# Benchmarking LLM Privacy Recognition for Social Robot Decision Making

### v1

@@ 5. LLM Evaluation
5.1. Model Selection & Setup
We evaluated five leading closed-source models, including GPT-o3 (o3-2025-04-16) (OpenAI, 2025b), GPT-4o (gpt-4o-2024-08-06) (OpenAI, 2025a), Gemini 2.5-Flash (gemini-2.5-flash-preview-04-17) (Google, 2025a), Gemini 2.5-Pro (gemini-2.5-pro-preview-05-06) (Google, 2025b), Grok 3 (grok-3-beta)(xAI, 2025), as well as five open-sourced models, including DeepSeek V3 (deepseek-chat) (DeepSeek, 2025), Qwen3-32B (Yang et al., 2025), Mistral-7B-Instruct-v0.3 (AI, 2024), Llama-3.3-70B-Instruct (Meta Platforms, 2024), Llama-4-Scout-17B-16E-Instruct (Meta Platforms, 2025).
As we assess each model, we utilize similar parameters; we set the temperature to 0.3 and the reasoning effort to ‘low’ for reasoning models. For each selected model, we run three trials to verify consistency in responses. If the overall agreement is greater than $90\%$, we accept the responses from the first trial. If the overall agreement is less than $90\%$ but greater than $70\%$, we accept the majority response for each item between the three trials. Two of the open-source models (i.e., Qwen and Mistral) are evaluated over two A600 48GB GPUs, and the remaining models are evaluated through their respective APIs (Huggingface Inference API and kluster.ai for the Llama series). For any given model or trial, we input questions to the LLMs in a JSON format.
@@ 5.2. Default Prompting Conformance
In this phase, we explore the privacy-related decision-making of out-of-box LLMs and the degree to which their responses conform to user expectations. To begin, we first compare each LLM’s responses with those of our participants. This process sets a baseline expectation for subsequent evaluation during which we provide the LLMs with elements of the user’s privacy attitudes and preferences. To properly compare the LLM responses with those of our participants, we present instructions to the LLMs that strictly resemble those provided during our user study. To that end, our model inputs are composed of a system prompt containing any necessary background information and output format, as well as a user prompt containing both scenarios and questions. A detailed example prompt is presented in Figure 6. Following data collection, we are left with responses to 15 survey questions for each of our 50 scenarios across each of the evaluated LLMs.
@@ 5.2.1. Analysis Procedure
To compare the data from a single LLM (i.e., one response per each of 50 scenarios) to the data of all participants (i.e., 27 responses per each of 50 scenarios), we expanded the LLM data to mirror our full participant data set. To do so, we simply duplicate the LLM data 27 times for ease of comparison across the 27 unique participants per scenario. Next, we compute the micro and macro F1 scores between the participant and LLM data, treating the human responses as the grounded truth. To compare numerical data (i.e., sensitivity rating), we calculate per-scenario mean absolute error (MAE).
@@
Llama Scout 0.487 0.868 0.209 0.531
GPT-o3 0.487 0.868 0.201 0.512
GPT-4o 0.487 0.868 0.197 0.499
Gemini 2.5 Pro 0.484 0.863 0.198 0.501
Llama 3.3-70B 0.487 0.868 0.191 0.481
DeepSeek V3 0.487 0.868 0.192 0.479
Grok3 0.487 0.868 0.190 0.476
Gemini 2.5 Flash 0.485 0.864 0.191 0.480
Qwen 3-32B 0.483 0.861 0.188 0.470
Mistral-7B-Instruct-v0.3 0.407 0.711 0.154 0.357
@@ 5.2.2. Results
We present our results in Tables 2 and 5. For comparison, we classify these results into three categories: binary (e.g., privacy-enhancing or non-interfering), categorical (i.e., three or more ordinal response options), and numerical (i.e., scaled from 0 to 10). Overall, performance between models is relatively consistent, with the exception of Mistral-7B-Instruct-v0.3, which performs notably worse among binary and categorical items. Across models, high micro F1 scores indicate that the models perform fairly well when compared to common user responses (i.e., when participants prefer privacy-enhancing behaviors). Low macro F1 scores, however, indicate that the models perform worse on rare responses (i.e., when participants prefer non-intrusive behaviors). While this pattern is present in both binary and categorical data, weaknesses in model performance become more pronounced in categorical data due to the increased number of possible responses. We also observe that eight models show a perfect tie among binary items, as these models selected privacy-enhancing choices for all items.
To further investigate why models prefer privacy-enhancing responses by default, we additionally examine the privacy orientation of the LLMs. As we see in Section 4, participant privacy orientation influences response preference. Therefore, we queried models to respond to the POS (Baruh and Cemalcılar, 2014) in the same manner we presented this scale to participants. The results of this process are shown in Figure 5. We see that nearly all models rate privacy as a right at the maximum (i.e., 5.00), with the exceptions Gemini 2.5 Flash, GPT-o3, and Mistral-7B-Instruct-v0.3 (4.67, 4.67, and 4.33, respectively). Similarly, nearly all models rate concern about privacy of others at the maximum, except Qwen 3-32B and Llama 4 Scout (4.80 and 4.40, respectively). Among the remaining two subscales, concern about own informational privacy and other-contingent privacy, which are less meaningful from the perspective of an LLM, all models scored between 3.50 and 4.50 and between 3.50 and 5.00, respectively. To the extent that these results can be interpreted as privacy orientations of the LLMs, they show that all models are highly concerned with the privacy of users.
Figure 5. LLM POS Scores—Privacy orientation scores for each of the evaluated LLMs. In cases where a user’s privacy may be concerned (i.e., Subscales 1 and 4), nearly all LLMs produce a maximal score.
@@
5.3. Model Conformity Under Prompting Strategies
@@
Model POS only Sensitivity only POS + Sensitivity Example + Sensitivity
@@
Llama Scout 0.487 0.858 0.181 0.506 0.516 0.643 0.250 0.492 0.540 0.759 0.224 0.511 0.603 0.746 0.373 0.592
o3 0.472 0.862 0.177 0.503 0.513 0.642 0.243 0.421 0.534 0.790 0.264 0.523 0.723 0.875 0.353 0.588
4o 0.493 0.858 0.175 0.504 0.517 0.642 0.244 0.438 0.503 0.839 0.193 0.489 0.664 0.810 0.392 0.640
Gemini 2.5 Pro 0.467 0.867 0.150 0.478 0.519 0.659 0.241 0.419 0.545 0.777 0.237 0.470 0.691 0.859 0.355 0.584
DeepSeek V3 0.506 0.850 0.177 0.468 0.522 0.672 0.211 0.396 0.532 0.821 0.204 0.463 0.692 0.837 0.399 0.612
Llama 3.3-70B 0.334 0.851 0.189 0.519 0.516 0.666 0.253 0.464 0.360 0.753 0.242 0.492 0.668 0.819 0.384 0.588
Grok3 0.485 0.854 0.185 0.506 0.519 0.650 0.239 0.430 0.525 0.825 0.231 0.513 0.719 0.885 0.397 0.628
Gemini 2.5 Flash 0.498 0.852 0.159 0.463 0.507 0.616 0.251 0.412 0.544 0.721 0.242 0.445 0.686 0.841 0.364 0.571
Qwen 3-32B 0.465 0.868 0.138 0.440 0.537 0.812 0.214 0.442 0.511 0.831 0.179 0.440 0.673 0.836 0.351 0.576
Mistral-7B-Instr. v0.3 0.494 0.820 0.135 0.339 0.494 0.851 0.149 0.447 0.490 0.764 0.139 0.377 0.231 0.862 0.335 0.582
Table 3. LLM Performance by Prompting Strategy—Conformity off LLMs to participant responses based on our four prompting strategies (i.e., POS prompting, sensitivity prompting, POS & sensitivity prompting, and few-shot prompting). These results include Micro and Macro F1 scores and are separated based on binary and categorical items.
@@
Sensitivity Difference 4.62% -16.53% 7.68% -4.25%
t-test $p=.000$ $p=.002$ $p=.000$ $p=.031$
POS & Sens. Difference 3.85% -6.25% 6.28% -0.63%
@@
Few-Shot Difference 16.53% -1.34% 21.77% 11.75%
@@ 5.3.1. POS Prompting
The first prompting strategy we use involves providing the LLMs with participant POS scores. As we have shown in Section 4.4.3, these scores, particularly subscale 1, are predictive of various preferences for robot responses. We therefore anticipate that the LLMs will be able to better align their responses to those of participants with this information. Rather than prompting the LLMs to evaluate each of the 50 scenarios only once, we instead request responses utilizing the POS scores of each participant across the specific scenarios each participant evaluated. As a result, we are able to obtain a dataset for each LLM that mirrors the dataset provided by participants. Following data collection, we again compute Micro and Macro F1 scores, separated by binary and categorical data, for each LLM.
@@ Results
Compared to default model performance, we see only modest and non-significant changes in conformity with human responses under this prompting strategy (see Table 4). As part of our analysis, we also examine how this prompting strategy influences LLM perceptions of scenario sensitivity. We do not observe a substantial improvement in this regard among the evaluated models, except for Qwen 3-32B (see Table 5). Moreover, several models even exhibit an increase in mean absolute error (MAE). The models that exhibit the smallest shift (see Figure 7), whether positive or negative, are the Llama series, the Gemini 2.5 series, and Grok3. These results suggest that even if user POS scores are provided to a model, they may mislead the model in terms of assessing the sensitivity of a given scenario. This possibility motivates our next prompting strategy, sensitivity prompting.
Model MAE $\downarrow$ Percent Reduction (%)
@@
Qwen 3-32B 3.311 2.575 $+$22.2
DeepSeek V3 3.155 2.827 $+$10.4
Grok 3 2.941 2.881 $+$2.1
Gemini 2.5 Flash 3.370 3.376 --–0.2
Llama 3.3-70B 2.774 2.794 --–0.7
Llama 4 Scout 2.685 2.778 --–3.4
Gemini 2.5 Pro 3.136 3.318 --–5.8
4o 2.773 3.053 --–10.1
Mistral-7B-Inst. 2.674 2.987 --–11.7
o3 2.824 3.241 --–14.8
@@ 5.3.2. Sensitivity Prompting
As participant and LLM sensitivity ratings vary widely, even when providing the LLMs with participant POS scores, we next attempt to provide the LLMs with participant sensitivity ratings directly for each participant-scenario pair. An example prompt is shown in Figure 6.
@@ 5.3.3. POS & Sensitivity Prompting
In an attempt to further improve model performance, we provide the LLMs with both participants’ POS scores and sensitivity ratings. Previously, POS prompting did not yield improved model performance and, in many cases, worsened model sensitivity ratings. Through sensitivity prompting, we saw increases in macro F1 scores and decreases in micro F1 scores. By providing both participant POS scores and sensitivity ratings (see Figure 6), we anticipate that the models will be able to utilize POS scores more effectively without hindering perceptions of scenario sensitivity.
@@ Results
While this prompting strategy performs better than sensitivity prompting among micro F1 scores, it performs worse in terms of macro F1 scores. Additionally, we only see a statistically significant change in micro F1 scores among binary questions and macro F1 scores among categorical questions compared to default prompting (see Figure 4). These results suggest that even when provided with user POS scores and sensitivity ratings, LLMs still have a limited understanding of human privacy expectations.
@@
5.3.4. Providing Examples and Sensitivity Scores
In all of the tests described above, the LLMs are limited to information regarding human perspectives, such as privacy orientation, concern for privacy in a given scenario. The LLMs are not, however, exposed to the actions participants would prefer in any given scenario. Therefore, our final prompting strategy is few-shot prompting (Brown et al., 2020). Under this strategy, we provide a comprehensive example of how a participant assesses a single scenario, including that participant’s POS scores, the sensitivity rating they provide, and their preferences regarding a robot’s responses. After providing this information to the LLMs, we then ask them to evaluate the remaining two scenarios that the same participant was provided. We include an example prompt in Figure 6.
Results
In Table 4, we see a statistically significant, large increase in macro F1 scores among binary questions as well as micro and macro F1 scores among categorical questions. While the micro F1 score for binary questions does not show a substantial nor statistically significant difference compared to default prompting. This result is expected, as both the default-prompted LLM and human responses primarily include privacy-preserving choices and already show high overall accuracy.
@@ 5.4. Case Study: LLM Logic & Reasoning
To further examine whether LLMs appropriately reason about participant privacy perspectives and preferences, we perform qualitative case studies on LLMs’ reasoning outputs.
@@ 5.4.1. Procedure
To begin, we selected our overall best- and worst-performing models, Grok3 and Mistral. For each of these models, we compared default and few-shot model results (see Table 2) along with the models’ rationales. For each comparison, we examined pairs of responses, based on the same scenario and participant responses for each, and analyzed the differences across four types of scenarios:
@@
    Both the default and few-shot results align perfectly with the participant responses.
@@
For the best-performing model, Grok3, we are able to identify all four scenarios defined above, whereas for the worst-performing model, Mistral, only cases (2) and (4) apply. Mistral never perfectly aligned with any participants (i.e., case 1) and never performed better under default prompting than few-shot prompting (i.e., case 3). Additionally, after restricting temperature, seeding, and top_k to be deterministic, Mistral would not generate responses matching those originally collected. Therefore, we add an additional step in prompting Mistral to include the original responses, and request that the model reason about the answers it previously made. In total, we produced 12 reasoning samples from these two models (i.e., 8 for Grok and 4 for Mistral). We next analyzed these cases to better understand the circumstances in which the models performed well or poorly.
5.4.2. Grok3
Overall, we observe that Grok3 provides a reasonable rationale for its responses, including providing references to details such as the POS scores, sensitivity ratings, and scenario-specific preferences of participants, as well as details about the scenarios themselves. Here are direct examples of Grok’s rationale relevant to its decision-making.

### v2

@@ 5.2. Default Prompting Conformance
In this phase, we explore the privacy-related decision making of out-of-the-box LLMs and the degree to which their responses conform to user expectations. To begin, we first compare each LLM’s responses with those of our participants. This process sets a baseline expectation for subsequent evaluation during which we provide the LLMs with elements of the users’ privacy attitudes and preferences. To properly compare the LLM responses with those of our participants, we present instructions to the LLMs that strictly resemble those provided during our user study. To that end, our model inputs are composed of a system prompt, containing any necessary background information and output format, as well as a user prompt, containing both scenarios and questions. A detailed example prompt is presented in Figure 6. Following data collection, we are left with responses to 15 survey questions for each of our 50 scenarios across each of the evaluated LLMs.
@@ 5.2.1. Analysis Procedure
To compare the data from a single LLM (i.e., one response per each of 50 scenarios) to the data of all participants (i.e., 27 responses per each of 50 scenarios), we expanded the LLM data to mirror our full participant data set. We duplicate the LLM data 27 times for ease of comparison across the 27 participants per scenario. Next, we compute Micro and Macro F1 scores between the participant and LLM data, treating the human responses as the ground truth. To compare numerical data (i.e., sensitivity rating), we calculate per-scenario mean absolute error (MAE).
@@
Llama 4 Scout 0.464 0.868 0.286 0.531
GPT-o3 0.464 0.868 0.270 0.512
Gemini 2.5-Pro 0.472 0.863 0.264 0.501
GPT-4o 0.464 0.868 0.231 0.499
DeepSeek V3 0.464 0.868 0.240 0.479
Gemini 2.5-Flash 0.469 0.864 0.236 0.480
Grok 3 0.464 0.868 0.229 0.476
Llama 3.3-70B 0.464 0.868 0.212 0.481
Qwen 3-32B 0.470 0.861 0.211 0.470
Mistral 7B-Instruct-v0.3 0.414 0.711 0.137 0.357
@@ 5.2.2. Results
We present our results in Tables 2 and 5. For comparison, we separate these results into three categories: binary (e.g., privacy-enhancing or non-interfering), categorical (i.e., three or more response options), and numerical (i.e., scaled from 1 to 10). Overall, performance between models is relatively consistent, except for Mistral, which performs notably worse among binary and categorical items. Across models, high Micro F1 scores indicate that the models perform fairly well when compared to common user responses (i.e., when participants prefer privacy-enhancing behaviors). Low Macro F1 scores, however, indicate that the models perform worse on rare responses (i.e., when participants prefer non-intrusive behaviors). While this pattern is present among all data, weaknesses in model performance become more pronounced in categorical data due to the increased number of possible responses. We also observe that six models show a perfect tie among binary items, as these models selected privacy-enhancing choices for all items.
To further investigate why models prefer privacy-enhancing responses by default, we additionally examine the privacy orientation of the LLMs. As shown in Section 4, participant privacy orientation influences response preferences. Therefore, we queried models to respond to the POS (Baruh and Cemalcılar, 2014) in the same manner we presented this scale to participants. The results of this process are presented in Figure 5. We see that nearly all models rate privacy as a right at the maximum (i.e., 5.00), except Gemini 2.5 Flash, GPT-o3, and Mistral (4.67, 4.67, and 4.33, respectively). Similarly, nearly all models rate concern about privacy of others at the maximum, except Qwen and Llama 4 Scout (4.80 and 4.40, respectively). Among the remaining two subscales, concern about own informational privacy and other-contingent privacy, which are less meaningful from the perspective of an LLM, all models scored between 3.50-4.50 and 3.50-5.00, respectively. To the extent that these results can be interpreted as privacy orientations of the LLMs, they indicate that all models are highly concerned with user privacy.
Figure 5. LLM POS Scores—Privacy orientation scores for each of the evaluated LLMs. Among subscales related to user privacy (i.e., 1 and 4), nearly all LLMs produce a maximal score.
@@
5.3. Conformity Under Prompting Strategies
@@
Model POS only Sensitivity only POS + Sensitivity Few-Shot
@@
Grok 3 0.484 0.854 0.280 0.506 0.514 0.650 0.319 0.430 0.516 0.825 0.332 0.513 0.706 0.885 0.476 0.628
GPT-o3 0.471 0.862 0.269 0.503 0.509 0.642 0.325 0.421 0.525 0.790 0.358 0.523 0.711 0.875 0.454 0.588
Llama 3.3-70B 0.481 0.851 0.277 0.519 0.512 0.666 0.327 0.464 0.519 0.753 0.312 0.492 0.658 0.819 0.458 0.588
GPT-4o 0.492 0.858 0.249 0.504 0.513 0.642 0.316 0.438 0.500 0.839 0.264 0.489 0.655 0.810 0.482 0.640
DeepSeek V3 0.505 0.850 0.262 0.468 0.518 0.672 0.274 0.396 0.525 0.821 0.285 0.463 0.682 0.837 0.485 0.612
Gemini 2.5-Pro 0.466 0.867 0.236 0.478 0.514 0.659 0.319 0.419 0.532 0.777 0.310 0.470 0.677 0.859 0.440 0.584
Llama 4 Scout 0.484 0.858 0.279 0.506 0.511 0.643 0.283 0.492 0.526 0.759 0.308 0.511 0.594 0.746 0.453 0.592
Qwen 3-32B 0.464 0.868 0.180 0.440 0.534 0.812 0.245 0.442 0.506 0.831 0.218 0.440 0.659 0.836 0.431 0.576
Gemini 2.5-Flash 0.496 0.852 0.208 0.463 0.502 0.616 0.307 0.412 0.534 0.721 0.285 0.445 0.671 0.841 0.446 0.571
Mistral 7B-Instr. v0.3 0.475 0.820 0.235 0.339 0.485 0.851 0.170 0.447 0.460 0.764 0.172 0.377 0.563 0.862 0.423 0.582
Table 3. LLM Performance by Prompting Strategy—Conformity of LLMs to participant responses based on our four prompting strategies (i.e., POS prompting, sensitivity prompting, POS & sensitivity prompting, and few-shot prompting). These results include Micro and Macro F1 scores and are separated based on binary and categorical items.
@@
Sensitivity Difference 5.01% -16.53% 5.67% -4.25%
@@
POS & Sens. Difference 5.32% -6.25% 5.28% -0.63%
[…]
Few-Shot Difference 19.67% -1.34% 22.33% 11.75%
@@ 5.3.1. POS Prompting
The first prompting strategy we use involves providing the LLMs with participant POS scores. As we have shown in Section 4.4.3, these scores, particularly Subscale 1, are predictive of various preferences for robot responses. We therefore anticipate that the LLMs will be able to better align their responses to those of participants with this information. Consequently, we generate responses that utilize the POS scores of each participant across the specific scenarios each participant evaluated. As a result, we obtain a dataset for each LLM that mirrors the dataset provided by participants. We again compute Micro and Macro F1 scores, separated by binary and categorical data, for each LLM.
@@ Results
Compared to default model performance, we primarily see modest and non-significant changes in conformity with human responses under this prompting strategy (see Table 4). As part of our analysis, we also examine how this prompting strategy influences LLM perceptions of scenario sensitivity. We do not observe a substantial improvement in this regard among the evaluated models, except for Qwen (see Table 5). Moreover, several models even exhibit an increase in mean absolute error (MAE). The models that exhibit the smallest shift (see Figure 7), whether positive or negative, are the Llama series, the Gemini 2.5 series, and Grok3. These results suggest that even if user POS scores are provided to a model, they may mislead the model in terms of assessing the sensitivity of a given scenario. This possibility motivates our next prompting strategy, sensitivity prompting.
Model MAE $\downarrow$
Percent
Reduction (%)
@@
Qwen 3-32B 3.311 2.575 ++22.2
DeepSeek V3 3.155 2.827 ++10.4
Grok 3 2.941 2.881 ++2.1
Gemini 2.5-Flash 3.370 3.376 -–0.2
Llama 3.3-70B 2.774 2.794 -–0.7
Llama 4 Scout 2.685 2.778 -–3.4
Gemini 2.5-Pro 3.136 3.318 -–5.8
GPT-4o 2.773 3.053 -–10.1
Mistral 7B-Inst. 2.674 2.987 -–11.7
GPT-o3 2.824 3.241 -–14.8
@@ 5.3.2. Sensitivity Prompting
As participant and LLM sensitivity ratings vary widely, even when providing the LLMs with participant POS scores, we next provide the LLMs with additional participant sensitivity ratings directly for each participant-scenario pair.
@@ 5.3.3. POS & Sensitivity Prompting
To further improve model performance, we provide the LLMs with both participants’ POS scores and sensitivity ratings. Previously, POS prompting did not yield improved model performance and, in many cases, worsened model sensitivity ratings. Through sensitivity prompting, we saw increases in Macro F1 scores and decreases in Micro F1 scores. By providing both participant POS scores and sensitivity ratings, we anticipate that the models will be able to utilize POS scores more effectively without hindering perceptions of scenario sensitivity.
@@ Results
While this prompting strategy performs better than sensitivity prompting among Micro F1 scores, it performs very similarly in terms of Macro F1 scores. Additionally, we see a statistically significant change in Micro and Macro F1 scores among binary questions and Macro F1 scores among categorical questions compared to default prompting (see Figure 4). These results suggest that even when provided with user POS scores and sensitivity ratings, LLMs still have a limited understanding of human privacy expectations.
5.3.4. Few-Shot Prompting
In all of the tests described above, the LLMs are limited to information regarding human perspectives, such as privacy orientation and concern for privacy in a given scenario. The LLMs are not, however, exposed to the actions participants would prefer in any given scenario. Therefore, our final prompting strategy is few-shot prompting (Brown et al., 2020). We provide a full example of how a participant assesses a single scenario, including that participant’s POS scores, the sensitivity rating they provide, and their preferences regarding a robot’s responses. After providing this information to the LLMs, we then ask them to evaluate the remaining two scenarios that the same participant was provided.
Results
In Table 4, we see a statistically significant, large increase in Macro F1 scores among binary questions as well as Micro and Macro F1 scores among categorical questions. The Micro F1 scores for binary questions do not show a substantial nor statistically significant difference compared to default prompting. This result is expected, as both the default-prompted LLM and human responses primarily include privacy-preserving choices and already show high overall accuracy.
@@ 5.4. Case Study: LLM Logic & Reasoning
To further examine whether LLMs appropriately reason about participant privacy perspectives and preferences, we performed qualitative case studies on our best-performing LLM’s reasoning outputs.
@@ 5.4.1. Procedure
To begin, we selected our overall best-performing model across all evaluation conditions, Grok3. We then compared Grok’s default and few-shot model results (see Table 2) along with its rationales. For each comparison, we examined pairs of responses, based on the same scenario and participant responses for each, and analyzed the differences across four types of scenarios:
@@
    Both the default and few-shot results align well with the participant responses.
@@
In total, we produced 32 reasoning samples (i.e., four pairs of samples across the four scenario types). We analyzed these cases to understand Grok’s performance in more depth.
5.4.2. Results
Overall, we observe that Grok provides a reasonable rationale for its responses, including providing references to details such as the POS scores, sensitivity ratings, and scenario-specific preferences of participants, as well as details about the scenarios themselves. Here, we describe direct examples of Grok’s rationale relevant to its decision making.